Find the final amount of money in an account if $6,800 is deposited at 4.5% interest compounded weekly and the money is left for six years. The final amount is? (round answer two decimal places.)
Answer by lwsshak3(11628)

Find the final amount of money in an account if $6,800 is deposited at 4.5% interest compounded weekly and the money is left for six years. The final amount is? (round answer two decimal places.)
**
Compound interest formula: A=P(1+i)^n, i=interest rate per compounding period, n=number of compounding periods, A= amt after n periods
For given problem:
n=6*52=312 (weeks)
i=.045/52
A=6800*(1+.045/52)^312
Using calculator:
A≈$8907 (amount of money after 6 years)